tyson:

The youth group is going on a trip to the state fair. The trip costs $58. Included in that price is $12 for a concert ticket and the cost of 2 passes, one for the rides and one for the game booths. Each of the passes costs the same price. Write an equation representing the cost of the trip, and determine the price of one pass. Solve your equation by showing your work and steps. (10 points)

Skatercat:

\[58-12\]

jordan5980:

2x + 12 = 58 <== ur equation (cost of the trip) 2x = 58 - 12 2x = 46 x = 46/2 x = 23 <== price of 1 pass

Skatercat:

i was still working on it T^T

tyson:

thx

jordan5980:

np